If a proton at a particular distance from a charged particle is repelled with a given force, by how much will the force decrease when the proton is 3 times farther away from the particle?


1
Expert's answer
2021-02-04T19:32:44-0500

The force between the proton and a charged particle can be written as follows:


"F=\dfrac{kq_1q_2}{r^2}."


If the proton is 3 times farther away from the particle the new force can be written as follows:


"F_{new}=\dfrac{kq_1q_2}{(3r)^2}=\dfrac{1}{9}\dfrac{kq_1q_2}{r^2}=\dfrac{1}{9}F."

Therefore, the force decreases to "\dfrac{1}{9}" its original value.